Support activity cost allocation

Jake’s Gems mines and produces diamonds, rubies, and other gems. The gems are produced by way of the Mining and Cutting activities. These production activities are supported by the Maintenance and Security activities. Security costs are allocated to the production activities based on asset value. Maintenance costs are normally allocated based on machine hours. However, Maintenance costs typically correlate more with the number of service calls. Information regarding the activities is provided in the following table:

Maintenance Security Mining Cutting
Number of service calls 17       20    60    20
Machine hours 89       88    182    176
Asset value $200,000       $80,000    $300,000    $300,000
Department cost $25,000       $42,500    $160,000    $95,000

1. Is the process of allocating maintenance costs based on machine hours correct? Identify the measure than can possibly be used to allocate the maintenance costs.

No

Measures that can possibly be used to allocate the maintenance costs.

a. Number of service calls

b. Asset value

c. Department cost

d. None of the above

a

2. Based on your response to part (1), determine the total costs allocated from each support activity to the other activities using the reciprocal services method and the most appropriate cost driver for Maintenance.

Maintenance
Department
Security
Department
Mining
Department
Cutting
Department
Maintenance cost allocation $ $ $
Security cost allocation $ $ $
Production activities total costs $ $

3. Assume Jake’s Gems is considering cutting costs by switching to a simpler support activity cost allocation method. Determine the total costs allocated from each support activity to the other activities.

Particulars Mining
Department
Cutting
Department
Maintenance cost $ $
Security cost $ $
Production activities total cost $ $
Answer #1

ANS 1) Jakes Gems mines had no the process of allocation maintenance cost on the basis of machine hours is incorrect. BecausWorking note: 1 Substituting the value of X in equation 2: Y = 42500 +0.20(25,000 +0.25Y) Y = 42500 +5,000+.05Y Y-0.05Y = 42,Reciprocal Allocation Method Number of Service calls % of Number of Service calls Asset Value % of Asset Value Department cos